# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v2

@@ III Our Contribution
The work described in this manuscript is the first-of-its-kind large-scale training of GFMs for atomistic materials modeling. We have employed three key techniques for developing a scalable and trustworthy GFM: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) ensemble methods for uncertainty quantification that allows model generalization and concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
@@ III-A Data Aggregation
Using large datasets for graph foundation model training can enhance generalizability and ensure resilience to data variance issues that typically arise during downstream tasks. To this end, we aggregated five open-source atomistic materials modeling datasets that are extremely diverse in terms of chemical composition, atomistic configurations, and number of atoms in the system. These datasets, as listed in Table I, are: ANI1x, QM7x, OC2020, OC2022, and MPTrj.
@@
  • OC2022 [45] provides 62,331 Density Functional Theory (DFT) relaxations (9,854,504 single point calculations) across a range of oxide materials, coverages, and adsorbates.
@@
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Fig. 1 shows the distribution of the number of atoms and bonds per molecule for each dataset. For the MPTrj dataset, approximately half of the molecules are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of molecules with different sizes and edge counts, with the larger molecules consisting of over 400 atoms and over 12,500 edges. In total, the data used for training our GFM consisted of 155 million molecules that consume 5.3 Terabytes of storage space. These datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O, as discussed in Section III-C.
Figure 1: The frequency distribution over number of atoms (upper) and edges (lower) for each dataset.
III-B Data Cleaning
Some of the atomistic structures were determined to have unrealistic values for atomic forces, in the order of 20,000 eV/angstrom. Thus, we first applied a data cleaning task in which we discarded all atomistic structures with an $L2$-norm (also known as spectral norm) of the force tensor above 100 eV/angstrom to ensure that these data samples did not affect the training of our GFMs. The number of data samples removed from each dataset by this filtering operation is reported in Table II.
@@ III-B Data Cleaning
TABLE II: Number of data samples discarded from each dataset in which the $L2$-norm (also known as spectral norm) of the force tensor was over 100 eV/angstrom.
@@ III-C Scalable Data Management
HydraGNN implements two optimization strategies that address scalability issues due to the large volume of data used for training. These strategies aim for 1) efficient storage and performant reading of large training data, and 2) fast reading of batch data during the training process. As molecular datasets are typically exported as collections of large numbers of files, storing datasets on a shared file system and then reading data from the large number of files during the training process causes a severe I/O bottleneck for GNN training. Multiple datasets cumulatively containing tens of thousands of small files put significant pressure on the filesystem’s metadata service, further slowing data access. Additionally, frequent data fetching by multiple GPUs from the file system during training loops results in a substantial slowdown in the training process. We adopted a two-pronged approach to managing large data and reducing the I/O overhead for training the GNN model. First, we pre-process the various input datasets and store their graph representation using a scientific data management library. Secondly, we use a distributed in-memory data store to load data into memory for fast shuffling of data objects during the training process.
@@ III-C1 ADIOS for High Performance I/O
Several publicly available molecular datasets are stored using bespoke schemas and exported as large collections of files. For example, the OC2020 dataset [44] consists of over 50,000 files. Storing multiple such datasets adds prohibitively high metadata overhead on the parallel file system and leads to slow data ingestion during the training process. For efficiently storing and performant reading of large training data, we use the ADIOS  [67] scientific data management library, which provides a state-of-the-art solution for managing extreme-scale data. ADIOS is designed to provide scalable I/O on the largest supercomputers in the world and has been successfully used in science applications that write and read several petabytes in a single simulation run.
@@
The number of sub-files controls the concurrency level while reading data in parallel. This $n:m$ pattern in which $n$ processes concurrently read data from $m$ sub-files is pivotal to obtaining high reading performance using ADIOS. ADIOS provides several options to tune I/O performance, including configuring the number of sub-files. We create the graph structures from input data and store them in ADIOS as a separate pre-processing step. We have developed a data writer and reader in HydraGNN for writing and reading graph data, respectively, from ADIOS files during the training process. When an ADIOS file is created, we split molecules into three groups - ‘trainset’ representing training data, ‘valset’ for data used for validation, and ‘testset’ data for testing the model performance. This logical grouping of molecules helps us read different groups of molecules for different tasks during the training process.
@@ III-C2 DDStore
Distributed data parallelism (DDP) [68, 69, 70, 71, 72] involves distributing training data amongst the available compute resources. Data is grouped into batches, and GPUs train on one batch at a time before fetching the next batch until all batches are processed in an epoch. Frequently reading data from the file system, even via a high-performance library such as ADIOS, is an expensive operation because I/O over the shared filesystem is the slowest operation in a computing system.
To provide fast data retrieval during training, we use DDStore [41], a distributed data store that provides in-memory data transfer between processes. When training begins, processes read data from ADIOS files and load into the node’s memory, which maintains a global map of data samples on each process. When a GPU requests a new batch of data, DDStore fetches the data from remote processes using low latency, fast communication techniques instead of reading data from the filesystem. By restricting access to the filesystem to the initial bootup phase, DDStore ensures that obtaining a batch is a fast, in-memory operation. Experiments described in [41] show that it leads to a 6$\times$ speedup in overall training time.
@@
III-D Scalable HPO
GNNs are known for their exceptional performance in learning from graph-structured molecular datasets. However, their development and broader application are hindered by the need for meticulous tuning of the network architecture. To achieve high predictive accuracy across chemically diverse datasets, it is essential to fine-tune the hyperparameters of HydraGNN. The task of identifying optimal hyperparameter settings is daunting and has been extensively documented in existing literature [73, 74, 75, 76, 77, 78]. Manual tuning requires extensive experimentation and often results in suboptimal performance.
@@ III-D Scalable HPO
Message passing is the core methodology of GNN models since it prescribes how features of nodes and edges are updated using information contained in neighboring nodes and edges. Various MPNNs have been developed and tailored for different atomistic systems, such as SchNet [80] for organic molecules and CGCNN [81] for solid state crystals. However, when considering foundation models applicable to a broad range of systems in atomistic materials, it is not practical to confine ourselves to a specific MPNN method. In HydraGNN, the choice of MPNN is configurable through a hyperparameter, allowing the users to select the optimal model that best suits their applications. We include MPNN as a categorical hyperparameter in the HPO runs to allow for the identification of the best performing MPNN layers for the assigned training data.
@@
In HPO, early termination strategies are vital for improving the utilization of computational resources by discarding unpromising candidates based on their performance trends. This decision has proven effective early in the training process [82]. Deephyper provides three early discarding techniques suited for asynchronous and parallel environments: (1) asynchronous successive halving, which progressively eliminates candidates based on their interim performance; (2) learning curve extrapolation, which predicts future performance from early data and facilitates early termination; and (3) constant fidelity, which sets a fixed resource allocation for each candidate before deciding whether to continue. For our tests, we used constant fidelity as it enables efficient reallocation of resources towards more promising configurations and significantly enhances operational efficiency in large-scale, distributed computing environments. We used 10 epochs as a stopping criterion for each model training in the HPO phase.
While HPO has been previously explored for GNNs, our approach uses HPO on a scale previously unattempted.
III-E Scalable Uncertainty Quantification with GNN Ensembles
@@
Ensemble methods are widely utilized in uncertainty quantification (UQ) to compile predictions from various models, termed ensemble members, into a unified forecast. The goal of these methods is to enhance model generalization by drawing on the diverse capabilities of each individual model.[83] To promote a varied set of predictions, practices such as different model initializations, techniques like Bagging and Boosting, and the integration of diverse network architectures are employed. Research conducted by Egele et al. [84] demonstrated that expanding the variety of network architectures within an ensemble can improve the diversity, thereby increasing the precision of uncertainty assessments. They also developed a technique for concurrently training multiple candidate models, which optimizes the use of computational resources. Ensemble methods are acknowledged for their ability to deliver reliable uncertainty estimates and their ease of implementation and scalability, making them practical for various UQ applications.
@@
Our method offers notable advantages in terms of generality and scalability. Central to our approach is the construction of model ensembles, which relies on scalable HPO. This methodology can be applied to any type of neural network model. The process begins with using a standard neural network architecture, conducting HPO, selecting the most suitable models, and subsequently producing uncertainty estimates. The scalability of our method is anchored in both the scalable nature of the hyperparameter search and the ability to train ensembles efficiently. Working with an ensemble of models enables many options for building consensus models, uncertainty estimation, and active learning.[83]
@@ IV-A Setup
We utilize three different sizes of foundation models for scaling measurements, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table III provides details about the three model sizes.
Experiments were conducted on two DOE supercomputers: Frontier at ORNL and Perlmutter at NERSC. Both systems provide state-of-the-art GPU-based heterogeneous architectures. Frontier, located at the Oak Ridge Leadership Computing Facility at ORNL, is one of the world’s most powerful supercomputers. It comprises a total of 9,408 compute nodes, each featuring a single 64-core AMD EPYC 7763 (Milan) CPU and four AMD MI250X GPU accelerators, effectively providing eight GPU units per node. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 64 GB of HBM2e (GPU) memory.
Perlmutter, a supercomputer at National Energy Research Scientific Computing Center (NERSC), features approximately 3000 CPU-only nodes and 1800 GPU-accelerated nodes. Our work utilizes only the GPU-accelerated nodes. Each node is equipped with an AMD EPYC 7763 (Milan) CPU and four NVIDIA Ampere A100 GPUs interconnected via NVLink-3. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 40 GB of HBM2 (GPU) memory. Both Frontier and Perlmutter use HPE Cray Slingshot(TM) interconnects.
To aid in monitoring HydraGNN execution in real-time for a subset of the analysis carried out on Frontier, an AMD research utility, omnistat, was employed to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, and frequency on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary prometheus server [85] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to one second.

### v3

@@ III Our Contribution
The work described in this manuscript is the first-of-its- kind large-scale training of GFMs for atomistic materials modeling using over 154 million atomistic structures as data samples and using over $91\%$ of the exascale supercomputer Frontier. We have employed three key techniques for developing a scalable and trustworthy GFM: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) ensemble methods for UQ that allows model generalization and concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
Compared to previous studies, our work shows near-linear scaling using 10x more GPUs and using much larger volumes of data, which introduces important challenges in I/O that we addressed to reduce computational bottlenecks and minimize communication overheads. Moreover, our results are generated using GPUs of newer generations, namely NVIDIA A100 installed on NERSC-Perlmutter and AMD Instinct™ MI250x installed on OLCF-Frontier, thereby showing that our scaling efficiency is also transferable across technologies manufactured by different vendors.
@@ III-A Data Aggregation
Using large datasets for GFM training can enhance generalizability and ensure resilience to data variance issues that typically arise during downstream tasks. To this end, we aggregated five open-source atomistic materials modeling datasets that are extremely diverse in terms of chemical composition, atomistic configurations, and number of atoms in the system. These datasets, as listed in Table I, are: ANI1x, QM7x, OC2020, OC2022, and MPTrj.
@@
  • OC2022 [44] provides 62,331 DFT relaxations (9,854,504 single point calculations) across a range of oxide materials, coverages, and adsorbates.
@@
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Figure 1 shows the distribution of the number of atoms and graph edges per atomistic structure for each dataset. For the MPTrj dataset, approximately half of the atomistic structures are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of atomistic structures with different sizes and edge counts, with the larger structures consisting of over 400 atoms and over 12,500 edges. In total, the data used for training, validating, and testing our GFM consisted of over 154 million atomistic structures that consume 5.3 Terabytes of storage space. These datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O, as discussed in Section III-C.
Figure 1: Normalized histograms for each dataset of the number of atoms within an atomistic structure (top) and number of edges within the graph representation of each atomistic structure (bottom).
Figure 2 provides the heatmap that illustrates the frequency of occurrence of each element of the periodic table across the entire dataset that results from the aggregation of the datasets ANI1x, QM7-X, OC2020, OC2022, and MPTrj. We notice that there is an under-representation of the elements from the transition metal groups. This is partially due to the fact that first-principles calculations for materials including these elements (e.g., alloys) are more computationally expensive to run, and thus are available in smaller volumes with respect to first-principles data for other classes of materials.
Figure 2: Heatmap that describes the frequency of occurrence of each element of the periodic table across data sampled resulting from the aggregation of the datasets ANI1x, QM7-X, OC2020, OC2022, and MPTrj.
III-B Data Cleaning and Pre-processing
Some of the atomistic structures were determined to have unrealistic values for atomic forces (on the order of 20,000 eV/angstrom). These probably corresponded to configurations visited at early stages of energy minimization during the generation of these datasets. To eliminate these outliers, we first applied a data cleaning operation in which we discarded all atomistic structures with an $L2$-norm (also known as spectral norm) of the force tensor above 100 eV/angstrom to ensure that these data samples did not affect the training of our GFMs. The number of data samples removed from each dataset by this filtering operation is reported in Table II.
@@ III-B Data Cleaning and Pre-processing
TABLE II: Number of data samples discarded from each dataset due to values of the $L2$-norm (also known as spectral norm) of the force tensor being unreasonably over 100 eV/angstrom.
The histogram of the values of the energy per atom and $L^{2}$-norm of forces, after removing data samples with force values unreasonably high are shown in Figure 3. The partition of each dataset into training, validation, and testing has been performed using the 80%-10%-10% splitting. In Figure 4 we show the histograms of the distribution of energies and atomic forces for the training, validation, and testing subsets of each dataset.
Figure 3: Normalized histograms for each dataset of the total energy per atom (top) and the $L^{2}$-norm of the force tensor (bottom) for each dataset after removing data samples with $L^{2}$-norm of the force tensor unreasonably higher than 100 eV/angstrom.
Figure 4: Histograms showing the energy and force distributions across five datasets, along with the distribution breakdown for the training, validation, and test sets. The y-axis shows the probability density on a logarithmic scale.
@@ III-C Scalable Data Management
HydraGNN implements two optimization strategies that address scalability issues due to the large volume of data used for training. These strategies aim for: 1) efficient storage and performant reading of large training data, and 2) fast reading of batch data during the training process. As atomistic materials modeling datasets are typically exported as collections of large numbers of files, storing datasets on a shared parallel file system (PFS) and then reading data from the large number of files during the training process causes a severe I/O bottleneck for GNN training. Multiple datasets cumulatively containing tens of thousands of small files put significant pressure on the PFS’s metadata service, further slowing data access. Additionally, frequent data fetching by multiple GPUs from the PFS during training loops results in a substantial slowdown in the training process. We adopted a two-pronged approach to manage large data and reducing the I/O overhead for training the GNN model. First, we pre-process the various input datasets and store their graph representation using a scientific data management library. Secondly, we use a distributed in-memory data store to load data into memory for fast shuffling of data samples during the training process.
@@ III-C1 ADIOS for High Performance I/O
Several publicly available atomistic materials modeling datasets are stored using bespoke schemas and exported as large collections of files. For example, the OC2020 dataset [43] consists of over 50,000 files. Storing multiple such datasets adds prohibitively high metadata overhead on the PFS and leads to slow data ingestion during the training process. For efficiently storing and reading of large training data, we use the ADIOS  [69] scientific data management library, which provides a state-of-the-art solution for managing extreme-scale data. ADIOS is designed to provide scalable I/O on the largest supercomputers in the world and has been successfully used in science applications that write and read several petabytes in a single simulation run.
@@
The number of sub-files controls the concurrency level while reading data in parallel. This $n:m$ pattern in which $n$ processes concurrently read data from $m$ sub-files is pivotal to obtaining high reading performance using ADIOS, which provides several options to tune I/O performance, including configuring the number of sub-files. We create the graph structures from input data and store them in ADIOS as a separate pre-processing step. We have developed a data writer and reader in HydraGNN for writing and reading graph data, respectively, from ADIOS files during the training process. When an ADIOS file is created, we split the data samples into three groups - ‘trainset’ representing training data, ‘valset’ for data used for validation, and ‘testset’ data for testing the model performance. This logical grouping of data samples helps us read different groups of data samples for different tasks during the training process.
@@ III-C2 DDStore
DDP [71, 72, 73, 74, 75] involves distributing training data amongst the available compute resources. Data is grouped into batches, and GPUs train on one batch at a time before fetching the next batch until all batches are processed in an epoch. Frequently reading data from the PFS, even via a high-performance library such as ADIOS, is an expensive operation because I/O over the shared PFS is the slowest operation in a computing system.
To provide fast data retrieval during training, we used DDStore [70], a distributed data store that provides in-memory data transfer between processes. When training begins, processes read data from ADIOS files and load into the node’s memory, which maintains a global map of data samples on each process. When a GPU requests a new batch of data, DDStore fetches the data from remote processes using low latency, fast communication techniques instead of reading data from the PFS. By restricting access to the PFS to the initial bootup phase, DDStore ensures that obtaining a batch is a fast, in-memory operation. Experiments described in [70] show that it leads to a 6$\times$ speedup in overall training time.
@@
III-D MTL for simultaneous prediction of energies and atomic forces
MTL uses a single DL architecture to simultaneously predict multiple quantities [76] and allows for a natural and automated incorporation of physics knowledge into the model by extracting correlations between the properties predicted, with manual intervention by a domain expert only needed in determining which quantities to use. The use of MTL is useful for developing GFMs, as it induces the model to use the physical correlations to develop physics-informed features that can be transferred on several downstream tasks. In this work, we choose energy and atomic forces as quantities to predict simultaneously. Denoting the number of atoms in an atomistic structure with $n$ and the Cartesian coordinates of the position of the nuclei of atom $i$ (where $i$ can be any integer between 1 and $n$) with $\mathbf{x}_{i}\in\mathbb{R}^{3}$, the relation between the energy $e\in\mathbb{R}$ and forces $\mathbf{f}_{i}\in\mathbb{R}^{3}$ acting on atom $i$ is
$\mathbf{f}_{i}=-\nabla_{\mathbf{x}_{i}}e.$ (1)
In HydraGNN, each predicted quantity is associated with a separate loss function and the global objective function minimized during the training is a linear combination of the individual loss functions. Denoting with $\mathbf{F}\in\mathbb{R}^{n\times 3}$ the atomic force tensor what contains atomic forces acting on all the $N$ atoms of a structure
$\mathbf{F}=\begin{bmatrix}\mathbb{f}^{T}_{1}\\ \vdots\\ \mathbb{f}^{T}_{N}\end{bmatrix},$ (2)
where $T$ represents the transposition of a vector. Denoting with $N$ the total number of parameters to train in the GFM architecture, the MTL loss function $\ell_{\mathrm{MTL}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{+}$ is:
$\ell_{\mathrm{MTL}}(\mathbf{W})=\alpha_{\mathrm{energy}}\lVert e_{\textrm{predict}}(\mathbf{W})-e\rVert_{1}+$ (3)
$\alpha_{\mathrm{forces}}\lVert\mathbf{F}_{\textrm{predict}}(\mathbf{W})-\mathbf{F}\rVert_{1},$ (4)
where $\lVert\cdot\rVert_{1}$ represents the $L_{1}$ norm of a vector or the induced tensor norm, $e$ and $\mathbf{F}$ represent the true values of energies and forces, $\mathbf{e}_{\textrm{predict}}$ and $\mathbf{F}_{\textrm{predict}}$ are the corresponding predictions given by HydraGNN, and the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ are used to calibrate the individual terms of the loss function for each individual properties. Similarly to [77], we calibrate the scalar positive weights $\alpha_{\mathrm{energy}}$ and $\alpha_{\mathrm{forces}}$ to account for the fact that the energy is a global quantity, while the atomic forces are local quantities. More specifically, we set the values of these scalar weights to $\alpha_{\mathrm{energy}}=1$ and $\alpha_{\mathrm{forces}}=100$. While it is helpful to optimize the weights as a hyperparameter, we found that this set of values is a suitable choice. An illustration of the main components of the HydraGNN architecture used for MTL is described in Figure 6.
Figure 6: HydraGNN architecture for simultaneous prediction of energy and atomic forces.
III-E Scalable HPO
GNNs are known for their exceptional performance in learning from graph-structured atomistic materials modeling datasets. However, their development and broader application are hindered by the need for meticulous tuning of the network architecture. To achieve high predictive accuracy across chemically diverse datasets, it is essential to fine-tune the hyperparameters of HydraGNN. The task of identifying optimal hyperparameter settings is daunting and has been extensively documented in existing literature [78, 79, 80, 81, 82, 83]. Manual tuning requires extensive experimentation and often results in suboptimal performance.
@@ III-E Scalable HPO
Message passing is the core methodology of GNN models since it prescribes how features of nodes and edges are updated using information contained in neighboring nodes and edges. Various MPNNs have been developed and tailored for different atomistic systems, such as SchNet [85] for organic molecules and CGCNN [86] for solid state crystals. However, when considering GFMs applicable to a broad range of systems in atomistic materials, it is not practical to confine ourselves to a specific MPNN method. In HydraGNN, the choice of MPNN is configurable through a hyperparameter, allowing the users to select the optimal model that best suits their applications. We include MPNN as a categorical hyperparameter in the HPO runs to allow for the identification of the best performing MPNN layers for the assigned training data.
@@
In HPO, early termination strategies are vital for improving the utilization of computational resources by discarding unpromising candidates based on their performance trends. This decision has proven effective early in the training process [87]. DeepHyper provides three early discarding techniques suited for asynchronous and parallel environments: (1) asynchronous successive halving, which progressively eliminates candidates based on their interim performance; (2) learning curve extrapolation, which predicts future performance from early data and facilitates early termination; and (3) constant fidelity, which sets a fixed resource allocation for each candidate before deciding whether to continue. For our tests, we used constant fidelity as it enables efficient reallocation of resources towards more promising configurations and significantly enhances operational efficiency in large-scale, distributed computing environments. We used 10 epochs as a stopping criterion for each model training in the HPO phase. While HPO has been previously explored for GNNs, our approach uses HPO on a scale previously unattempted.
III-F Scalable UQ with GNN Ensembles
@@
Ensemble methods are widely utilized in UQ to compile predictions from various models, termed ensemble members, into a unified forecast. The goal of these methods is to enhance model generalization by drawing on the diverse capabilities of each individual model [88]. To promote a varied set of predictions, practices such as different model initializations, techniques like Bagging and Boosting, and the integration of diverse network architectures are employed. Research conducted by Egele et al. [89] demonstrated that expanding the variety of network architectures within an ensemble can improve the diversity, thereby increasing the precision of uncertainty assessments. They also developed a technique for concurrently training multiple candidate models, which optimizes the use of computational resources. Ensemble methods are acknowledged for their ability to deliver reliable uncertainty estimates and their ease of implementation and scalability, making them practical for various UQ applications.
@@
Our method offers notable advantages in terms of generality and scalability. Central to our approach is the construction of model ensembles, which relies on scalable HPO. This methodology can be applied to any type of NN model. The process begins with using a standard NN architecture, conducting HPO, selecting the most suitable models, and subsequently producing uncertainty estimates. The scalability of our method is anchored in both the scalable nature of the hyperparameter search and the ability to train ensembles efficiently. Working with an ensemble of models enables many options for building consensus models, uncertainty estimation, and active learning [88].
@@ IV-A Setup
For strong and weak scaling studies we utilize three different sizes of GFM architectures, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table III provides details about the three model sizes.
Experiments were conducted on two DOE supercomputers: Frontier at ORNL and Perlmutter at NERSC. Both systems provide state-of-the-art GPU-based heterogeneous architectures. Frontier, located at the Oak Ridge Leadership Computing Facility at ORNL, is currently the world’s fastest and most powerful supercomputer [90]. It comprises a total of 9,408 compute nodes, each featuring a single 64-core AMD EPYC 7763 (Milan) CPU and four AMD Instinct MI250X GPU accelerators, effectively providing eight GPU units per node. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 64 GB of HBM2e (GPU) memory.
Perlmutter, a supercomputer at National Energy Research Scientific Computing Center (NERSC), features approximately 3000 CPU-only nodes and 1800 GPU-accelerated nodes. Our work utilizes only the GPU-accelerated nodes. Each node is equipped with an AMD EPYC 7763 CPU and four NVIDIA Ampere A100 GPUs interconnected via NVLink-3. Running with one rank per GPU unit, each rank has 64 GB of DDR4 (CPU) and 40 GB of HBM2 (GPU) memory. Both Frontier and Perlmutter use HPE Cray Slingshot(TM) interconnects.
To aid in monitoring HydraGNN execution in real-time for a subset of the analysis carried out on Frontier, an AMD Research utility, Omnistat [91], was employed to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, energy consumption, and clock/memory frequencies on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a Prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary Prometheus server [92] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to one second.